OpenStudy (sugarapples):

at a town a large convention center, the cost of a hotel room has increased by 5.1% annually. if the average hotel room cost $ 48.00 in 1980 and this growth continues, what will the average hotel room cost in 2012 use this formula y=a91+r)^t and round to the nearest cent a. $143.38 b. $126.34 c. $235.79 d. $87.19 please help i will fan and give a medal

OpenStudy (anonymous):

d

OpenStudy (anonymous):

can u help me with a question since I have helped you?

Nnesha (nnesha):

do you the formula ??

OpenStudy (sugarapples):

y=a(1+r)^t

Nnesha (nnesha):

a= ?? r= ?? t = ?? there is a key word "annually" what does that mean ??

Nnesha (nnesha):

nvm formula is already up there hmm so a= ? r= ? t = ?

OpenStudy (anonymous):

The formula to find the perimeter of a regular polygon is P = n × s. This means that you multiply the number of sides times the length of each side. Which polygon will have the greatest perimeter? A. a triangle with each side 7 inches B. a square with each side 4 inches C. an octagon with each side 2 inches D. a pentagon with each side 5 inches

Nnesha (nnesha):

@home_work welcome to open study ?? please post your question at LHS where it says ":ask a queston...."

OpenStudy (anonymous):

post does not work

OpenStudy (sugarapples):

a= 5.1 r= 2012 t= 48 i think that is how it goes

Nnesha (nnesha):

r = 5.1% rate u need it in decimal so divide by 100 t = time a = initial amount (cost) "average hotel room cost $ 48.00 in 1980" but you need average in 2012 right so subtract 1980 -2012 = ?

Nnesha (nnesha):

so a= ?? and t =?? r = ??

OpenStudy (anonymous):

do u know my question

OpenStudy (sugarapples):

for the subtracting it would be 32

Nnesha (nnesha):

okay what would you substitute for a and for r ? t = year which is 32

OpenStudy (sugarapples):

5.1 and 48

Nnesha (nnesha):

5.1% change that into decimal

OpenStudy (sugarapples):

0.051

Nnesha (nnesha):

yep right a =48 t=32 r= .051 now replace variables by their values \[\huge\rm y =a(1+r)^{t}\]

OpenStudy (sugarapples):